OpenStudy (anonymous):

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete this job on his own

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

x(x-8)/(x+x-8) = 3 solve for x

OpenStudy (michele_laino):

the working rate of Jimmy is W/x, where W is the work to be done. The working rate of Harry is W(x-8) whereas the working rate when both Jimmy and Harry work together is: \[\frac{W}{x} + \frac{W}{{x - 8}}\]

OpenStudy (anonymous):

I understand:)

OpenStudy (michele_laino):

so, using your data we can write: \[\Large \frac{W}{x} + \frac{W}{{x - 8}} = \frac{W}{3}\] or after a simplification: \[\Large \frac{1}{x} + \frac{1}{{x - 8}} = \frac{1}{3}\] please solve that last equation for x

OpenStudy (michele_laino):

oops..the working rate of Harry is W/x-8

OpenStudy (anonymous):

So now do we do?

OpenStudy (anonymous):

so we now find a common denominator?

OpenStudy (michele_laino):

the common denominator, is: 3x(x-8)

OpenStudy (michele_laino):

and we got this equivalent equation: \[\Large 3\left( {x - 8} \right) + 3x = x\left( {x - 8} \right)\] with the condition: x-8>0 since x-8 is a time which has to be positive

OpenStudy (anonymous):

ok, ok:)

OpenStudy (michele_laino):

I simplify that equation, nd I get this: \[\Large \begin{gathered} 3x - 24 + 3x = {x^2} - 8x \hfill \\ {x^2} - 14x + 24 = 0 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

uh huh! :)

OpenStudy (michele_laino):

the solution of that quadratic equation are: \[\Large \begin{gathered} x = \frac{{14 + \sqrt {{{\left( { - 14} \right)}^2} - 4 \times 1 \times 24} }}{{2 \times 1}} = 12 \hfill \\ \hfill \\ x = \frac{{14 - \sqrt {{{\left( { - 14} \right)}^2} - 4 \times 1 \times 24} }}{{2 \times 1}} = 2 \hfill \\ \end{gathered} \] please check my values. Only x012 is acceptable

OpenStudy (michele_laino):

oops..only x=12 is acceptable

OpenStudy (anonymous):

so the first equation works?

OpenStudy (anonymous):

Thanks:) So Jimmy can work 12 hours alone?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

yes! Jimmy works for x=12 hours only

OpenStudy (anonymous):

Thanks!

OpenStudy (michele_laino):

:)